OpenStudy (keana):

What is the product of m^2-m-6/8p * 2p/2m+m^2? a: m+3/4m b: m-3/4m c: m+3/4m

OpenStudy (anonymous):

Can you factorize it? \[m^2 - m - 6\]

OpenStudy (keana):

@waterineyes i don't know

OpenStudy (anonymous):

Then I think you should reply fast if you don't know..

OpenStudy (keana):

@waterineyes i had to do something

OpenStudy (anonymous):

\[m^2 -m - 6\] Here find two factors of 6 such that their product is -6 and their sum is -1: -3 and 2.. -3*2 = -6 -3 + 2 = -1 Both condition satisfy: So: -3 and 2 are the factors: Now: \[m^2 -3m + 2m - 6 \implies m(m-3) + 2(m-3) \implies (m-3)(m+2)\]

OpenStudy (anonymous):

What can you factor out of : \(2m + m^2\) ??

OpenStudy (keana):

@waterineyes i still do not know !

OpenStudy (anonymous):

What you don't know ??

OpenStudy (keana):

@waterineyes what you can factor !!

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

In 2m and \(m^2\) what is common ??

OpenStudy (anonymous):

(psst. its the m)

OpenStudy (keana):

m @waterineyes

OpenStudy (anonymous):

So by factor out it means taking m common or outside like: \[m^2 + 2m\] \[m(m+2)\] This is called taking a factor out..

OpenStudy (keana):

@waterineyes okk ..

OpenStudy (anonymous):

So now your question has become: \[\frac{(m-3)(m+2)}{8p} \times \frac{2p}{m(m+2)}\] Now your job is to cancel out the things which are same in numerator and denominator..